To help prevent frost damage, 4kg of 0°c water is sprayed unto a fruit tree.
How much heat transfer occurs as the water freezes?
How much would the Temperature of the 120kg tree decrease if this amount of heat is transferred from the tree assume that no face change occur(SHC of the tree as 3.35j/kj/k)
1
Expert's answer
2019-07-29T08:18:53-0400
"Q_{freeze}=mL=4(334)=1336 \ kJ"

"Q_{freeze}=Mc\Delta T"

"\Delta T=\frac{Q_{freeze}}{Mc}=\frac{1336}{(120)(3.35)}=3.32\ K"